OpenStudy (babynini):

write the expression as a sum or difference and simplify cos6m sin (-4m)

OpenStudy (babynini):

@optiquest ...me again :}

OpenStudy (babynini):

\[\cos \alpha \sin \beta = 1/2[\sin(\alpha+\beta) - \sin (\alpha-\beta)]\]

OpenStudy (babynini):

= 1/2[sin(6m+(-4m)) - sin(6m - (-4m))]

OpenStudy (babynini):

= 1/2 [ sin (2m) - sin (10m)]

OpenStudy (babynini):

Right? How do I simplify it further?

OpenStudy (babynini):

@zepdrix ... :}

OpenStudy (babynini):

can it be 1/2[-sin8] ? Can I get rid of the half somehow?

OpenStudy (babynini):

= -sin4 ??

zepdrix (zepdrix):

Oh you again -_- buhahaha jk. sec i gotta read the stuff

OpenStudy (babynini):

I sowriiee heh just 7 more questions after this! xP D:

OpenStudy (babynini):

#sleepdeprivation

zepdrix (zepdrix):

quote bebnini: = 1/2 [ sin (2m) - sin (10m)] end quote. perfect!

OpenStudy (babynini):

cannot be simplified further? ooo

zepdrix (zepdrix):

`write the expression as a sum or difference` and simplify you did this. good. i dunno, simplify is sooooo subjective. trigonometry is very circular, you could simplify for days, turning it into other things. don't try to overdo it. this is a good place to stop.

OpenStudy (babynini):

that's so true D: the prof is super particular and loves things to be simplified to the veery simplest. Which is why I always am making sure. But ok :)

zepdrix (zepdrix):

ya, so much ambiguity in the word simplify when you're dealing with trig D: ugh, sucks to have a professor like that lol

OpenStudy (babynini):

for reals. but it's not illegal to simplify it extra like I did? or you dont' know?